OpenStudy (anonymous):

What is the slope of the line that passes through the points (1, 2) and (-2, 4)?

OpenStudy (lalaly):

\[\underline{slope}= \frac{ \Delta y}{ \Delta x}\]

OpenStudy (anonymous):

can you explain how i would get the slope?

OpenStudy (diyadiya):

\[Slope = \frac{y_2-y_1}{x_2-x_1}\]\[\frac{4-2}{-2-1}=\frac{-2}{3}\]

OpenStudy (anonymous):

thank you